MATEMATIQKI VESNIK

59 (2007), 151

UDK 517.518.5 originalni nauqni rad

research paper

Corrigendum to

SPLINE-WAVELET SOLUTION OF SINGULARLY PERTURBED BOUNDARY PROBLEM

Desanka Radunović

Due to a technical error, Table 3 in the paper “Spline-wavelet solution of singularly perturbed boundary problem” by Desanka Radunović, published in the last issue of Matematčki Vesnik 59, 1–2 (2007), 31–46, appeared wrongly. The right Table 3 should be

\begin{matrix}ε\h & 2^{−3} & 2^{−4} & 2^{−5} & 2^{−6} & 2^{−7} & 2^{−8} \\ 1 & 6e − 4 & 2e − 4 & 4e − 5 & 1e − 5 & 2e − 6 & 5e − 7 \\ 0.1 & 1e − 1 & 3e − 2 & 8e − 3 & 2e − 3 & 5e − 4 & 1e − 4 \\ 0.01 & 3e + 0 & 1e + 0 & 4e − 1 & 1e − 1 & 3e − 2 & 7e − 3 \\ 0.001 & 1e + 2 & 2e + 1 & 7e + 0 & 3e + 0 & 5e − 1 & 4e − 2 \\ 0.0001 & 1e + 4 & 2e + 3 & 2e + 2 & 4e + 1 & 2e + 1 & 7e + 0 \\ 0.00001 & 1e + 6 & 1e + 5 & 2e + 4 & 3e + 3 & 4e + 2 & 1e + 2 \end{matrix}

Table 3. Uniform errors for problem (22) in quadratic case

We apologize to the author and the readers.